please help me with 3,4,5
Thank you!








3. Which sample had more moles of HA in it initially? Give reas on ! 4. Do the prelab calculation found in part IIIA below 5. Calculate the (expected) initial concentration of your acid solution (see Part IIIB below to get the initial volume of your acid solution) Acid 2 0 50 40 30 0 10 20 Volume of titrant (base) added (mL) III. Experimental Information and Procedures A.Background and preliminary calculation. You will be given access to one of the following four acids: Mandelic Acid (CoH,CH(OH)COOH; 152.16 g/mol); (ii) Glycolic Acid (CH2(OH)COOH; 76.05 g/mol); (lifi) KHC H,O4 (an acidic salt, 204.23 g/mol); (iv) NaHPO HO (an acidic salt, 138.0 g/mol) You will be preparing a solution containing your acid, and then you will titrate that solution with a standardized solution of approximately 0.1 M NaOH ("standardized" means that its concentration is known with great accuracy and precision; record the value from the NaOH(aq) reagent bottle). Prelab Calculation: Given the molar mass of your acid (just pick one), calculate the number of grams of the acid needed to just react with 30.0 mL of 0.100 M NaOH. Now, given what you just calculated, if you were to start with that number of grams of acid in a titration, at what "volume of titrant added" would you expect to reach the equivalence point if the titrant used were 0.100 M NaOH? (Note this is NOT a trick question, I just want to make sure you understand what the meaning of "equivalence point" is in a titration!) I hape that you realize that the answer to this question is "30.0 mL" Please ask me if you don't understand this. B. Preparation of the acid solution. 1. TARE a clean and dry 100-mL beaker on the balance (i.e., put the beaker on the pan and press the TARE button; the reading should go to 0.000 g), and then weigh out the number of grams of your acid that you calculated in part A into the beaker (to within 0.01 g of what you calculated; do time trying to get it to match more precisely than that). waste 2. Add 40.0 mL of deionized water using a graduated cylinder, and swirl to dissolve all of the solid. 3. Add a magnetic stir bar, and possibly a few drops of an acid-base indicator (optional; ask me).
14 I. Prelab (Practice in interpreting titration curves) 1. For each titration curve, find (be precise in a & bl): a. The equivalence point volume and pH b. The half equivalence point volume and pH. c. The K, of the acid being titrated. 2. Which acid is the weaker acid? How do you know? 3. Which sample had more moles of HA in it initially? Give reason! 4. Do the prelab calculation found in part IIA below. 5. Calculate the (expected) initial concentration of your acid solution (see Part IlIB below to get the initial volume of your acid solution). 12 10 Acid 1 Acid 2 60 50 40 20 30 10 Volume of ttrant (base) added imi) II. Exnerimental Informationcs
Answer #1

3. Knowing that both acids were titrated with the same base, we could say that there was more acid2 moles at the beggining because it takes more volume base to get to equivalence point

4.

NaOH mole = (Volume) * (Molarity) = (30 x10-3 L) * ( 0.1 mol/L) = 3 x10-3 mole

to react with the amount of NaOH now known, we will use the same of amount of Acid, this means

NaOH mole = Acid mole

and now we can pick one of the four acids to calculate the grams with

Acid grams = (Acid mole) * (Molecular weight of the acid)

for example we can take, Glycolic Acid

Glycolic Acd grams = (3 x10-3 mole) * (76.05 g/mole) = 228.15 x10-3 gram = 0.22815 gram

Now we can say that equivalence will be reached when 30 mL of NaOH were added due to there will be the same amount of mole, Acid mole = Base mole

5. The expected concentration will be given by Molarity formula, with these values

V = 40 mL = 40 x10-3 L

grams = Acid grams = 0.22815 gram (Calculated previously)

Molecular weight of the acid = MW = 76.05 g/mole (As used previously)

M = Molarity

M= grams/(MW*V)

M = (0.22815 grams)/(76.05 g/mole * 40 x10-3 L) = 0.075 M